OpenStudy (campayne):

Use a graph and a table to solve the system. Check your answer. Enter your answer as the following example: (5, -3). {3y + 6x = 3 x - y = -7

OpenStudy (kittiwitti1):

Here, they want you to graph the two equations given, and see if they reach the point (\(5,-3\)) at all. I would put both equations in point-slope form, \(y=mx+b\), so it is easier to graph them. For instance, \(x-y=-7\) becomes \(y=x-(-7)\) or \(y=x+7\).

OpenStudy (kittiwitti1):

Let's try \(3y+6x=3\). How would you put that in the form of \(y=mx+b\)?

OpenStudy (campayne):

umm y=-2x+1

OpenStudy (campayne):

right

OpenStudy (kittiwitti1):

Yeah. Sorry

OpenStudy (kittiwitti1):

So now you have the equations:\[y=x+7\]\[y=-2x+1\]All you need to do now is graph those and then see if they both run into the point (\(5,-3\)) somewhere.

OpenStudy (campayne):

its not looking like it does

OpenStudy (kittiwitti1):

Okay, then your answer is No

OpenStudy (campayne):

but thats not what its asking

OpenStudy (kittiwitti1):

Oops, I did the first equation wrong.\[x-y=-7\]\[y=-x-7\]

OpenStudy (kittiwitti1):

... I'm not exactly sure what they want for "solve the system" sorry

OpenStudy (campayne):

its all good all good you at least helped me with SOMETHING lol

OpenStudy (kittiwitti1):

True, true xD